OpenStudy (anonymous):

Perform the indecation. (36m^-60m+25) divided by (6m-5= 6m+5, 6m-5 R. 50, 6m+15R. 100, or 6m-5?

OpenStudy (anonymous):

here you go :) (36m^2 - 60m + 25) ÷ (6m - 5) =[(6m)^2 - 2*6m*5 + 5^2] ÷ (6m - 5) =(6m - 5)^2 ÷ (6m - 5) =(6m - 5)

OpenStudy (anonymous):

does this make sense??

OpenStudy (anonymous):

let me see :)

OpenStudy (anonymous):

lol do you need help with another one??

OpenStudy (anonymous):

yes if u can :) I'm doing an exam and have 48 to do.

OpenStudy (anonymous):

okey dokey whats the next one?

OpenStudy (anonymous):

can i write down what u showed me first?

OpenStudy (anonymous):

kk here we go!

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

Select the correct product. (x2^+6x+9)(3x-1) what would be the right one for this?

OpenStudy (anonymous):

thanks you soo much for your help :)

OpenStudy (anonymous):

-6p^3(3p^2+5p-1 whats the ans?

OpenStudy (anonymous):

The correct product of (x2 6x 9)(3x - 1) is 54x^3(3x-1).

OpenStudy (anonymous):

for the second one the answer is -18p^5 - 30p^4 + 6p^3

OpenStudy (anonymous):

wait the (x2^+6x+9)(3x-1) well the answer is 3x^3 + 17x^2 + 21x - 9

OpenStudy (anonymous):

now that makes sence now:)

OpenStudy (anonymous):

lol anymore?

OpenStudy (anonymous):

yeah lots more lol. let me put down the answer and then i'll be bach

OpenStudy (anonymous):

okey dokey what are you in Algebra?

OpenStudy (anonymous):

grade or unit? 9th grade on the last unit

OpenStudy (anonymous):

(x+8)(y+3

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

factor 16x^2+24+9.

OpenStudy (anonymous):

u?

OpenStudy (anonymous):

for the first it is: xy + 3x + 8y + 24<<< answer

OpenStudy (anonymous):

hah your pic looks like a girl from a band that I've seen live :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

(4x + 3)^2 here's for the second one

OpenStudy (anonymous):

and lol that's weird

OpenStudy (anonymous):

no it's cool if u ask me.

OpenStudy (anonymous):

another one

OpenStudy (anonymous):

yea I guess so lol whats your next question?

OpenStudy (anonymous):

factor the four- term polynomial. pq-2r+pr-2q.

OpenStudy (anonymous):

The band is Barlow Girl ( christian rock chick band)

OpenStudy (anonymous):

(p - 2)(q + r) lol here you go and oh wow I love that chick ! I love the song by them called never alone

OpenStudy (anonymous):

factor the following polynomial completely. -x2^y2^+x4^+4y2^-4x2^

OpenStudy (anonymous):

(x + 2)(x - 2)(x + y)(x - y) there you go :)

OpenStudy (anonymous):

omg!!! u know them too? yeah people say I look like Lauren ( the drummer) U look like Becca

OpenStudy (anonymous):

haha that's awesome

OpenStudy (anonymous):

more.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

the scientific notation 2.15x10-3^ has what value? sry the _ between the 10 and 3 is supposed to be the same high as the 3 so in the middle of those numbers but higher up.

OpenStudy (anonymous):

2,150

OpenStudy (anonymous):

answers: 0.000215, 0.00215, 215, or 2,150.

OpenStudy (anonymous):

lol its 2,150

OpenStudy (anonymous):

sry i messed up here a spaceship traveled 93 billion miles. In scientific notation that is _______ miles?

OpenStudy (anonymous):

find the product. (y3^)2^.y7^

OpenStudy (anonymous):

for the first one : 9.3 x 10^10

OpenStudy (anonymous):

given the sequence 2,6,18,54......, which expression shown would give the fifteen term?

OpenStudy (anonymous):

here you go: we would use the rule that if (a^m)^n = a^mn and also that (a^m)*(a^n) = a^(m+n) so in this case we got (y^6)*y^7 and now we use the second rule because it has the same variable to obtain y^(6+7) = y^13 :)

OpenStudy (anonymous):

another one

OpenStudy (anonymous):

here's for the second one T(n) = 2/3 x 3n T(15) = 2/3 x 315

OpenStudy (anonymous):



OpenStudy (anonymous):